OpenStudy (anonymous):

the length of a rectangle is 5 cm less than twice its width. the perimeter of the rectangle is 38 cm. what are the dimensions of the rectangle ?

OpenStudy (anonymous):

l=2w-5 2l+2w=38 set up a system of eqns then plug in l 2(2w-5)+2w=38 4w-10+2w=38 6w-10=38 6w=48 w=8 width is 8 cm l=2(8)-5 l=11cm

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (anonymous):

np :) hope u understood how i got that

OpenStudy (anonymous):

yes i did! very well :D